OpenStudy (anonymous):

Grace is making party hats for her upcoming birthday party. A single hat has a diameter of 21 centimeters and a slant height of 38 centimeters. What is the lateral area of each party hat? 1,599.86 cm2 3,892.43 cm2 2,506.99 cm2 1,253.50 cm2

Parth (parthkohli):

To start with, a party hat has a shape of a cone.

Parth (parthkohli):

Also, the lateral area would be \({1\over2}\pi d\times \text{slant height}\)

OpenStudy (anonymous):

@AlexPR787 @bobbylynn @RioAnne @DannyRomero19 @april115 @marcoduuuh

OpenStudy (mathmate):

Are you going to try ParthKohli's formula?

OpenStudy (anonymous):

my answer is b

Parth (parthkohli):

Nope.

OpenStudy (anonymous):

y not

OpenStudy (mathmate):

Evaluate \[ Area\ =\ \frac{1}{2}\pi d (slant\ height) \] and you'd find out.

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

0.5

OpenStudy (mathmate):

and...

OpenStudy (mathmate):

d=21 cm slant height = 38 cm, so area = ?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

0.2763

OpenStudy (anonymous):

Try this on a calulator: 0.5 * 3.14 * 21 * 38

Parth (parthkohli):

That's what I input in Google. ^

OpenStudy (anonymous):

really^^^

OpenStudy (anonymous):

1252.86

OpenStudy (mathmate):

Hurray! If you had used 3.1416, you would have got the exact value.

OpenStudy (anonymous):

my answer is the 1 one

OpenStudy (mathmate):

Would you like to look at all the answers before deciding? :)

OpenStudy (anonymous):

i did

OpenStudy (mathmate):

Which one is the closest to 1252?

OpenStudy (anonymous):

last one

OpenStudy (mathmate):

I agree!

OpenStudy (anonymous):

thnk you so much

OpenStudy (mathmate):

You're welcome! :)